The human ear can hear sounds at frequencies from 20 Hz- 20,000 Hz. What are the wavelength equivalent to the lowest frequency a human can hear? What will the size of the organ pipe that will play in it is wavelength? The pipeline is 1/2 of the wavelength.
1
Expert's answer
2019-11-06T09:23:39-0500

We have

"\nu_{min}=20Hz"

Then the wavelength equivalent to it

"\lambda_{max}=\frac{v_{sounds}}{\nu_{min}} = \frac{331\text{ m/s}}{20\text{ 1/s}}=16.55m"

organ pipe size is

"L=\frac{1}{2} \cdot \lambda_{max} =\frac{1}{2} \cdot 16.55=8.275m"